I need help dolving an equation?
a=3,c=3, B=64 degrees

Answer by stanbon(75887)

a=3,c=3, B=64 degrees
----
Law of Cosines to find "b":
b^2 = a^2+c^2-2ac*cos(64)
------
b^2 = 9 + 9 - 2*3*3*0.4384
---
b^2 = 18 - 7.8907
----
b^2 = 10.1093
b = 3.1795
-----------------
If a = c the angles opposite are Equal:
A + 64 + C = 180 degrees
----
A+C = 116 degrees
A = 116/2 = 58 degrees = C
===============================
Cheers,
Stan H.
==============